Use the information below to answer questions 33-3
Answer #1

we know that

concentration = moles / volume (L)

so

initially

[HI] = 0.66 / 2 = 0.33

now

2HI ----> H2 + I2

using ICE table

at equilibrium

[HI] = 0.33 -2x

[H2] = x

[I2] = x

now

Kc = [H2] [I2] /[HI]^2

so

0.0156 = [x] [x] / [0.33-2x]^2

0.0156 = [x]^2 / [0.33-2x]^2

solving we get

x = 0.033

so

33)

at equilibrium

[H2]= x = 0.033 M

option E) 0.033 M

34)

[I2] = x = 0.033 M

option E) 0.033 M

35)

[HI] = 0.33 - 2x

[HI] = 0.33 - ( 2 * 0.033)

[HI] = 0.264 M

option A) 0.264 M